# SLIM-Brain: A Data- and Training-Efficient Foundation Model for fMRI Data Analysis

### v1

@@ A.4 Internal experiments
We fine-tune our model and evaluate it on the held-out $20\%$ split of the HCP dataset (sex classification and subject identification/fingerprinting) as summarized in Table 5. We benchmark against two representative fMRI foundation models, BrainMass (yang2024brainmass), and Brain-JEPA (dong2024brain)—using the authors’ released checkpoints (pretrained on 3232k–6464k samples). For a controlled comparison under matched data budgets, we additionally retrain each baseline on the same total number of sessions as ours and report results for their best validation checkpoints (Table 5 and Fig. 1).
@@ A.5 Linear probing Results
To assess the intrinsic quality of the learned representations—independent of task-specific adaptation—we conduct linear probing (LP). LP freezes the pretrained encoder and trains a lightweight linear classifier on top, providing a direct measure of how well the latent features capture task-relevant information and their linear separability he2022masked. We evaluate LP on two downstream tasks, focusing on ADNI diagnosis (Table 6). Despite pretraining on a small subset of data, our linear probes transfer robustly to out-of-distribution (OOD) settings, outperforming baseline encoders and, in some cases, even their fine-tuned counterparts. At the same time, compared to MAE-based models, our method shows less performance degradation when using linear probing (2.5% vs. 5.1%).

### v2

@@ A.4 Internal experiments
We fine-tune our model and evaluate it on the held-out $20\%$ split of the HCP dataset (sex classification and subject identification/fingerprinting) as summarized in Table 5. We benchmark against two representative fMRI foundation models, BrainMass (Yang et al., 2024), and Brain-JEPA (Dong et al., 2024)—using the authors’ released checkpoints (pretrained on 3232k–6464k samples). For a controlled comparison under matched data budgets, we additionally retrain each baseline on the same total number of sessions as ours and report results for their best validation checkpoints (Table 5 and Fig. 1).
@@ A.5 Linear probing Results
To assess the intrinsic quality of the learned representations—independent of task-specific adaptation—we conduct linear probing (LP). LP freezes the pretrained encoder and trains a lightweight linear classifier on top, providing a direct measure of how well the latent features capture task-relevant information and their linear separability He et al. (2022). We evaluate LP on two downstream tasks, focusing on ADNI diagnosis (Table 6). Despite pretraining on a small subset of data, our linear probes transfer robustly to out-of-distribution (OOD) settings, outperforming baseline encoders and, in some cases, even their fine-tuned counterparts. At the same time, compared to MAE-based models, our method shows less performance degradation when using linear probing (2.5% vs. 5.1%).